# arXiv:2602.23790v1[cs.CV] 27 Feb 2026

### 27 Feb 2026 v1

@@ 1 Introduction
Recent progress in ROD focuses on designing rotation-sensitive convolutions [32, 38, 51], modifying the backbone [22, 1, 45] or detection neck [36, 29, 27, 35, 23], or optimizing angle regression losses [37, 14, 43, 44]. In this work, we propose Fourier Angle Alignment (FAA) that leverages frequency domain to estimate and align object orientation to enhance existing ROD methods. As shown in Fig.1, our plug-and-play modules can be applied on the main-stream ROD architectures to deal with two common bottlenecks: directional incoherence in the neck and task conflict in the detection head.
When giving features from different levels, as shown in Fig.2, traditional FPN [25] fuses them using simple element-wise addition or concatenation. However, on the one hand, features from high levels commonly have strong semantics but blurred spatial details, thus making the orientation coarse and low-frequency [4]. On the other hand, features from low levels have sharp edges and textures, so they carry precise, high-frequency direction cues. When these two types of features are directly fused, their orientation signals often conflict. This mismatch creates noise in the fused features, degrading angle prediction accuracy.
Since the same RoI feature must serve two different tasks, classification and angle regression often conflict. As shown in Fig.3, classification needs rotation-invariant features, for a plane should always be recognized as a plane regardless of its orientation. Regression, however, needs rotation-sensitive features, where the feature must change when the object rotates. Forcing a single feature to satisfy both goals leads to a compromise: the feature becomes neither fully invariant nor fully sensitive. This limits the model’s ability to accurately predict both class and angle.
@@ 2.2 Frequency-domain Analysis
Frequency-domain analysis has a long and rich history in computer vision. Early approaches [40, 48, 28] primarily used Fourier properties to design rotation-invariant handcrafted feature extractors, then used these features for downstream tasks. In recent years, with the rise of deep learning, researchers [21, 47, 9, 8, 49] have begun to explicitly integrate Fourier transforms into neural network architectures to exploit their global receptive fields and inherent properties under geometric transformations. For instance, FcaNet [33] replaces global average pooling with a frequency-channel attention mechanism, modeling long-range dependencies in the spectral domain to significantly boost performance in image classification and object detection. More recently, a series of works have underscored the critical role of explicit frequency modeling in dense prediction. Spatial Frequency Modulation [6] mitigates aliasing during downsampling via a modulation-demodulation scheme to preserve high-frequency details. FreqFusion [4] decomposes features into low- and high-frequency components to jointly enhance intra-class consistency and boundary sharpness. Frequency-Adaptive Dilated Convolution [5] dynamically adjusts dilation rates and kernel spectral responses based on local frequency content to balance receptive field size and effective bandwidth. And modules like FDAM [7] counteract the low-pass bias of Vision Transformers through high-frequency attention compensation. Collectively, these advances demonstrate that explicit spectral modeling has established a powerful foundation for improving the performance of feature extraction, thus providing a new view of oriented object detection.
@@ 3.1 Formulation and Motivation
Oriented Object Detection.
Let the input image be $\textbf{I}(\textbf{x})$, $\textbf{x}=(x,y)^{\intercal}$. For a horizontal object detector $\mathcal{D}_{\text{HBB}}$, the output is a set of axis-aligned bounding boxes with class information:
@@
Fourier Rotation Equivariance.
Let $\textbf{I}(\textbf{x})$ denote a signal in the spatial domain. Its frequent spectrum is
@@
which means if a spatial signal is rotated by $\phi$, its frequent spectrum will also be rotated by $\phi$, as shown in Fig 4.
Spectral alignment of rectangles.
Suppose $\textbf{I}_{\text{rec}}(\mathbf{x})$ denote a 2D rectangular function in the spatial domain and $\textbf{x}=(x,y)^{\intercal}$ Specifically,
@@
Since $a>b$, the main lobe of $\operatorname{sinc}(2au)$ is narrower than that of $\operatorname{sinc}(2bv)$. Hence, spectral energy decays more slowly along the $v$-axis than along the $u$-axis. Consequently, in the high-frequency regime, the dominant spectral energy lies along the $v$-axis, which is perpendicular to the major axis of the rectangle, as shown in Fig 4.
Motivations.
These observations reveal a fundamental property: object orientation can be reliably estimated from the frequency domain. This offers a direct and principled way to handle orientation, which is largely ignored by current detectors that rely only on spatial features.
@@ 3.2 Fourier Angle Alignment
We propose Fourier Angle Alignment (FAA), with the main progress shown in Fig.5. Generally speaking, FAA can be divided into two steps: Fourier Angle Estimation and Angle Alignment.
Fourier Angle Estimation.
With a square feature map $\mathbf{X}\in\mathbb{R}^{H\times H}$ as input, the FAA estimates its dominant orientation and performs rotational alignment. To begin with, FAA applis a 2D discrete Fourier transform to the input feature map to obtain its spectral representation:
@@
Angle Alignment.
After obtaining the estimated main direction, FAA rotates the original feature map by a specific reference angle to align the main direction with this reference angle:
@@ 5 Conclusion
In this work, we propose Fourier Angle Alignment, leveraging fourier rotation equivariance and spectral alignment of rectangles to better model orientation cues in oriented object detection tasks. We theoretically show that the spectral energy of a rectangular object concentrates perpendicular to its major axis, and use this insight to estimate dominant orientations in the frequency domain and align features accordingly. We integrate FAA into both neck and the detecting head, designing FAAFusion and FAA Head, which enhance orientation consistency across scales and improve rotation invariance of RoI features with minimal computational overhead. Experiments on three benchmark datasets demonstrate the efficacy of our approach in remote sensing object detection. In the future, it is worth exploring the generalization of FAA to other vision tasks such as instance segmentation or remote sensing changing detection, and investigate more efficient frequency-domain alignment strategies for real-time applications.

### 13 Mar 2026 v2

# Fourier Angle Alignment for Oriented Object Detection in Remote Sensing

@@ 1 Introduction
Recent progress in ROD focuses on designing rotation-sensitive convolutions [35, 41, 54, 56], modifying the backbone [25, 1, 48] or detection neck [39, 32, 30, 38, 26], or optimizing angle regression losses [40, 15, 46, 47]. In this work, we propose Fourier Angle Alignment (FAA) that leverages frequency domain to estimate and align object orientation to enhance existing ROD methods. As shown in Figure 1, our plug-and-play modules can be applied on the main-stream ROD architectures to deal with two common bottlenecks: directional incoherence in the neck and task conflict in the detection head.
When giving features from different levels, as shown in Figure 2, traditional FPN [28] fuses them using simple element-wise addition or concatenation. However, on the one hand, features from high levels commonly have strong semantics but blurred spatial details, thus making the orientation coarse and low-frequency [4]. On the other hand, features from low levels have sharp edges and textures, so they carry precise, high-frequency direction cues. When these two types of features are directly fused, their orientation signals often conflict. This mismatch creates noise in the fused features, degrading angle prediction accuracy.
Since the same RoI feature must serve two different tasks, classification and angle regression often conflict. As shown in Figure 3, classification needs rotation-invariant features, for a plane should always be recognized as a plane regardless of its orientation. Regression, however, needs rotation-sensitive features, where the feature must change when the object rotates. Forcing a single feature to satisfy both goals leads to a compromise: the feature becomes neither fully invariant nor fully sensitive. This limits the model’s ability to accurately predict both class and angle.
@@ 2.2 Frequency-domain Analysis
Frequency-domain analysis has a long and rich history in computer vision. Early approaches [43, 51, 31, 22] primarily used Fourier properties to design rotation-invariant handcrafted feature extractors, then used these features for downstream tasks. In recent years, with the rise of deep learning, researchers [24, 50, 10, 52, 23] have begun to explicitly integrate Fourier transforms into neural network architectures to exploit their global receptive fields and inherent properties under geometric transformations. For instance, FcaNet [36] replaces global average pooling with a frequency-channel attention mechanism, modeling long-range dependencies in the spectral domain to significantly boost performance in image classification and object detection. More recently, a series of works[8, 9] have underscored the critical role of explicit frequency modeling in dense prediction. Spatial Frequency Modulation [6] mitigates aliasing during downsampling via a modulation-demodulation scheme to preserve high-frequency details. FreqFusion [4] decomposes features into low- and high-frequency components to jointly enhance intra-class consistency and boundary sharpness. Frequency-Adaptive Dilated Convolution [5] dynamically adjusts dilation rates and kernel spectral responses based on local frequency content to balance receptive field size and effective bandwidth. And modules like FDAM [7] counteract the low-pass bias of Vision Transformers through high-frequency attention compensation. Collectively, these advances demonstrate that explicit spectral modeling has established a powerful foundation for improving the performance of feature extraction, thus providing a new view of oriented object detection.
@@ 3.1 Formulation and Motivation
Oriented Object Detection. Let the input image be $\textbf{I}(\textbf{x})$, $\textbf{x}=(x,y)^{\intercal}$. For a horizontal object detector $\mathcal{D}_{\text{HBB}}$, the output is a set of axis-aligned bounding boxes with class information:
@@
Fourier Rotation Equivariance. Let $\textbf{I}(\textbf{x})$ denote a signal in the spatial domain. Its frequent spectrum is
@@
which means if a spatial signal is rotated by $\phi$, its frequent spectrum will also be rotated by $\phi$, as shown in Figure 4.
Spectral alignment of rectangles. Suppose $\textbf{I}_{\text{rec}}(\mathbf{x})$ denote a 2D rectangular function in the spatial domain and $\textbf{x}=(x,y)^{\intercal}$ Specifically,
@@
Since $a>b$, the main lobe of $\operatorname{sinc}(2au)$ is narrower than that of $\operatorname{sinc}(2bv)$. Hence, spectral energy decays more slowly along the $v$-axis than along the $u$-axis. Consequently, in the high-frequency regime, the dominant spectral energy lies along the $v$-axis, which is perpendicular to the major axis of the rectangle, as shown in Figure 4.
Motivations. These observations reveal a fundamental property: object orientation can be reliably estimated from the frequency domain. This offers a direct and principled way to handle orientation, which is largely ignored by current detectors that rely only on spatial features.
@@ 3.2 Fourier Angle Alignment
We propose Fourier Angle Alignment (FAA), with the main progress shown in Figure 5. Generally speaking, FAA can be divided into two steps: Fourier Angle Estimation and Angle Alignment.
Fourier Angle Estimation. With a square feature map $\mathbf{X}\in\mathbb{R}^{H\times H}$ as input, the FAA estimates its dominant orientation and performs rotational alignment. To begin with, FAA applis a 2D discrete Fourier transform to the input feature map to obtain its spectral representation:
@@
Angle Alignment. After obtaining the estimated main direction, FAA rotates the original feature map by a specific reference angle to align the main direction with this reference angle:
@@ 5 Conclusion
In this work, we propose Fourier Angle Alignment (FAA), leveraging fourier rotation equivariance and spectral alignment of rectangles to better model orientation cues in oriented object detection tasks. We theoretically show that the spectral energy of a rectangular object concentrates perpendicular to its major axis, and use this insight to estimate dominant orientations in the frequency domain and align features accordingly. We integrate FAA into both neck and the detecting head, designing FAAFusion and FAA Head, which enhance orientation consistency across scales and improve rotation invariance of RoI features with minimal computational overhead. Experiments on three benchmark datasets demonstrate the efficacy of our approach in remote sensing object detection. In the future, it is worth exploring the generalization of FAA to other vision tasks such as instance segmentation or remote sensing changing detection, and investigate more efficient frequency-domain alignment strategies for real-time applications.
Acknowledgement
This work is supported by the National Natural Science Foundation of China (62331006), and the Fundamental Research Funds for the Central Universities.